# Hemispheric-Specific Coupling Improves Modeling of Functional Connectivity Using Wilson–Cowan Dynamics

### v1

@@ 2.1.   Empirical Dataset and Brain Network Estimation
We employed a publicly available dataset containing diffusion and functional MRI scans from 27 patients with schizophrenia and 27 demographically matched healthy controls [2, 5]. These data were acquired using a 3T Siemens Trio scanner and include high-resolution T1-weighted images, diffusion spectrum imaging (DSI), and resting-state fMRI. Participants in the patient group met DSM-IV criteria for schizophrenia or schizoaffective disorder, and were recruited from the Lausanne University Hospital. Brain network construction was performed using the Connectome Mapping Toolkit [6]. Cortical segmentation was based on the Lausanne 2008 multi-scale parcellation scheme [7], from which we selected the 83-region resolution to balance anatomical specificity with computational feasibility. SC matrices were derived from deterministic streamline tractography applied to the DSI data, with edge weights defined as normalized streamline counts accounting for surface area and length biases [8]. FC was estimated as the absolute Pearson correlation between regional mean BOLD signals after standard preprocessing, including motion correction, nuisance regression, spatial smoothing, and band-pass filtering [9, 10]. This framework provides subject-level structural and functional networks with matched parcellations, enabling direct comparison and simulation of brain dynamics constrained by individual connectomes. To reduce inter-subject variability and highlight group-level trends, we computed group-average SC and FC matrices separately for the control (CTRL) and schizophrenia (SCHZ) cohorts. Since all individual SC and FC matrices are symmetric by construction—owing to the undirected nature of tractography and the use of Pearson correlation—the resulting group-average matrices are also symmetric. This approach allows us to focus on global network properties and systematic differences between groups, rather than individual-specific fluctuations. By simulating neural dynamics on the group-averaged SC using the Wilson–Cowan model, we can directly assess how alterations in anatomical connectivity influence the emergence of functional interactions at the population level. Furthermore, the use of averaged matrices facilitates a controlled exploration of model parameters, such as global versus hemispheric-specific coupling schemes, enabling us to investigate how different modes of interaction (intra-hemispheric vs. inter-hemispheric connectivity) affect the model’s ability to reproduce empirical FC. This is particularly relevant in schizophrenia, where disruptions in inter-hemispheric integration have been consistently reported. Finally, from a translational perspective, the group-level SC and FC matrices serve as clinical reference profiles, against which deviations at the individual level could be later evaluated. This provides a meaningful scaffold for assessing the impact of pathological alterations and testing hypotheses about network-level mechanisms underlying psychiatric disorders. The 83-region resolution includes both cortical and subcortical structures. To restrict the analysis to cortical dynamics, we excluded subcortical regions and the brainstem, resulting in 68 cortical ROIs (34 per hemisphere). This reduction ensures compatibility with the Wilson–Cowan neural mass model, which is defined at the mesoscopic cortical level. All modeling and simulation steps were performed on these $68\times 68$ group-average matrices.
2.2.   Wilson–Cowan Network Model
@@ 3.1.   Topological Analysis of Empirical Structural Connectivity
In Fig. (2a) displays the histogram of SC weights based on group-average matrices. Both groups exhibit a predominance of weak connections, consistent with the known sparsity of long-range anatomical projections. However, the SCHZ group shows a slightly higher count of weaker edges, suggesting a reduction in overall anatomical connectivity strength compared to the CTRL group. In Fig. (2b), the histogram of node betweenness centrality reveals that the CTRL group tends to distribute intermediary structural roles across a broader range of nodes, as indicated by a slightly wider spread in centrality values. In contrast, the SCHZ group exhibits a more concentrated profile, with fewer nodes contributing to long-range communication pathways. Nonetheless, the overall difference between the two groups remains modest at the level of group-average SC. In Fig. (2c), the histogram of node degree suggests that patients with schizophrenia exhibit a more homogeneous structural connectivity in terms of degree, with fewer extremes (hubs or peripheral nodes), which may imply alterations in the network’s capacity for global integration and efficient communication.
@@ 3.2.   Wilson–Cowan Model with Uniform Global Coupling $G$
Figure 3: Best-performing simulation for the SCHZ group using global coupling $G=1.0$. (a) Excitatory activity, (b) Hilbert envelope, (c) SC, (d) Simulated FC (e) and empirical FC. The simulation yielded $r=0.37$, RMSE $=0.22$.
@@
Fig. (4) shows the average Pearson correlation ($\braket{r}$) and the average root mean square error ($\braket{\text{RMSE}}$) across simulations as a function of the global coupling parameter $G$, along with one standard deviation. The CTRL group reached a peak $\braket{r}=0.27\pm 0.03$, while the SCHZ group achieved a higher maximum of $\braket{r}=0.30\pm 0.03$, both groups at $G=1.0$. In both cases, the RMSE curve reached its minimum near the same values of $G$, reinforcing the identification of optimal coupling regimes. Specifically, the lowest RMSE values were $\braket{\text{RMSE}}=0.25\pm 0.01$ for the CTRL group and $\braket{\text{RMSE}}=0.20\pm 0.01$ for the SCHZ group, both groups occurring at $G=1.0$. On average, the SCHZ group achieved higher correlation values across simulations, suggesting that the model reproduces functional connectivity more accurately in patients than in controls under this global coupling scheme.
3.3.   Wilson–Cowan Model with Distinct Coupling Parameters $G_{1}$ and $G_{2}$
@@
Figure 5: Best-performing simulation for the SCHZ group using hemispheric-specific coupling parameters $G_{1}=0.80$, $G_{2}=15.00$. (a) Structural Connectivity (SC) matrix, (b) Simulated Functional Connectivity (FC), and (c) Empirical FC. The simulation yielded a correlation of $r=0.47$ and RMSE = 0.20.
@@ 4.   Conclusions
In this study, we implemented the Wilson–Cowan model and introduced a hemispheric-specific coupling scheme. We examined both groups and compared structural connectivity network features. In structural connectivity, we found results consistent with previous studies reporting decreased anatomical connectivity strength [13, 14], redistribution of central roles across network regions [14], and topological reorganization in schizophrenia [1]. To further explore the mechanistic basis of these alterations, we evaluated the capacity of a biologically grounded neural mass model—based on Wilson–Cowan dynamics [3, 4, 12], in order to reproduce resting-state functional connectivity from empirical cortical structural connectomes in matched control and schizophrenia groups. In our implementation, we differentiated intra-hemispheric and inter-hemispheric anatomical projections via a hemispheric-specific coupling scheme. This modified version, referred to as our model throughout the paper, systematically outperformed the classic global coupling approach, yielding higher Pearson correlations between simulated and empirical FC across both groups. The improvement was especially pronounced in the schizophrenia group in both models. This may reflect a stronger constraint of functional dynamics by the underlying structural scaffold in schizophrenia, where reduced global integration and increased modularity limit the repertoire of functional configurations. In such conditions, anatomically grounded models may more accurately capture the structure–function mapping [15, 11]. Interestingly, the optimal parameter configuration differed between groups. The SCHZ group achieved a higher maximum correlation of $\langle r\rangle=0.43\pm 0.03$ at $G_{1}=0.8$, $G_{2}=15.0$, whereas the CTRL group reached $\langle r\rangle=0.36\pm 0.05$ at $G_{1}=1.0$, $G_{2}=15.0$. Although both groups showed optimal performance in regimes where inter-hemispheric coupling $G_{2}$ exceeded intra-hemispheric coupling $G_{1}$, the schizophrenia group required a lower $G_{1}$ to reproduce empirical FC. This may suggest a diminished role of intra-hemispheric integration in shaping functional dynamics in schizophrenia, potentially reflecting degraded or less functionally effective intra-hemispheric anatomical pathways [14]. Additionally, the SCHZ model displayed good performance across a broader range of $G_{2}$ values, indicating a degree of flexibility or compensatory reliance on inter-hemispheric interactions. Together, these findings suggest distinct structure–function coupling regimes between groups, with schizophrenia exhibiting both reduced intra-hemispheric integration and a more tolerant configuration space for inter-hemispheric coupling. Overall, these findings highlight the value of integrating structural connectivity data with biologically inspired models to gain mechanistic insights into the altered brain dynamics observed in neuropsychiatric disorders.

### v2

@@ 2.1.  Empirical Dataset and Brain Network Estimation
We employed a publicly available dataset containing diffusion and functional MRI scans from 27 patients with schizophrenia and 27 demographically matched healthy controls [13, 14]. These data were acquired using a 3T Siemens Trio scanner and include high-resolution T1-weighted images, diffusion spectrum imaging (DSI), and resting-state fMRI. Participants in the patient group met DSM-IV criteria for schizophrenia or schizoaffective disorder, and were recruited from the Lausanne University Hospital. Brain network construction was performed using the Connectome Mapping Toolkit [6]. Cortical segmentation was based on the Lausanne 2008 multi-scale parcellation scheme [7], from which we selected the 83-region resolution to balance anatomical specificity with computational feasibility. SC matrices were derived from deterministic streamline tractography applied to the DSI data, with edge weights defined as normalized streamline counts accounting for surface area and length biases [3]. FC was estimated as the absolute Pearson correlation between regional mean BOLD signals after standard preprocessing, including motion correction, nuisance regression, spatial smoothing, and band-pass filtering [11, 4]. This framework provides subject-level structural and functional networks with matched parcellations, enabling direct comparison and simulation of brain dynamics constrained by individual connectomes. To reduce inter-subject variability and highlight group-level trends, we computed group-average SC and FC matrices separately for the control (CTRL) and schizophrenia (SCHZ) cohorts.Since all individual SC and FC matrices are symmetric by construction, owing to the undirected nature of tractography and the use of Pearson correlation, the resulting group-average matrices are also symmetric. This approach allows us to focus on global network properties and systematic differences between groups, rather than individual-specific fluctuations. By simulating neural dynamics on the group-averaged SC using the Wilson–Cowan model, we can directly assess how alterations in anatomical connectivity influence the emergence of functional interactions at the population level. Furthermore, the use of averaged matrices facilitates a controlled exploration of model parameters, such as global versus hemispheric-specific coupling schemes, enabling us to investigate how different modes of interaction (intra-hemispheric vs. inter-hemispheric connectivity) affect the model’s ability to reproduce empirical FC. This is particularly relevant in schizophrenia, where disruptions in inter-hemispheric integration have been consistently reported. Finally, from a translational perspective, the group-level SC and FC matrices serve as clinical reference profiles, against which deviations at the individual level could be later evaluated. This provides a meaningful scaffold for assessing the impact of pathological alterations and testing hypotheses about network-level mechanisms underlying psychiatric disorders. The 83-region resolution includes both cortical and subcortical structures. To restrict the analysis to cortical dynamics, we excluded subcortical regions and the brainstem, resulting in 68 cortical ROIs (34 per hemisphere). This reduction ensures compatibility with the Wilson–Cowan neural mass model, which is defined at the mesoscopic cortical level. All modeling and simulation steps were performed on these $68\times 68$ group-average matrices.
$2.$  Wilson–Cowan Network Model
@@ 3.1.  Topological Analysis of Empirical Structural Connectivity
Fig. (2a) displays the histogram of SC weights based on group-average matrices. Both groups exhibit a predominance of weak connections, consistent with the known sparsity of long-range anatomical projections. However, the SCHZ group shows a slightly higher count of weaker edges, suggesting a reduction in overall anatomical connectivity strength compared to the CTRL group. In Fig. (2b), the histogram of node betweenness centrality reveals that the CTRL group tends to distribute intermediary structural roles across a broader range of nodes, as indicated by a slightly wider spread in centrality values. In contrast, the SCHZ group exhibits a more concentrated profile, with fewer nodes contributing to long-range communication pathways. Nonetheless, the overall difference between the two groups remains modest at the level of group-average SC. In Fig. (2c), CTRL and SCHZ degree distributions are largely similar, with only subtle shifts, though the SCHZ group appears to exhibit fewer high-degree hubs.
@@ 3.2.  Wilson–Cowan Model with Uniform Global Coupling $G$
Fig. (4) shows the average Pearson correlation ($\braket{r}$) and the average root mean square error ($\braket{\text{RMSE}}$) across simulations as a function of the global coupling parameter $G$, along with one standard deviation. The CTRL group reached a peak $\braket{r}=0.27\pm 0.03$, while the SCHZ group achieved a higher maximum of $\braket{r}=0.30\pm 0.03$, both groups at $G=1.0$.
Figure 3: Group-averaged best-performing simulation for the SCHZ group using global coupling $G=1.0$. (a) Excitatory activity, (b) Hilbert envelope, (c) SC, (d) Simulated FC (e) and empirical FC. The simulation yielded $r=0.37$, RMSE =0.22=0.22.
@@
In both cases, the RMSE curve reached its minimum near the same values of $G$, reinforcing the identification of optimal coupling regimes. Specifically, the lowest RMSE values were $\braket{\text{RMSE}}=0.25\pm 0.01$ for the CTRL group and $\braket{\text{RMSE}}=0.20\pm 0.01$ for the SCHZ group, both groups occurring at $G=1.0$. On average, the SCHZ group achieved higher correlation values across simulations, suggesting that the model reproduces functional connectivity more accurately in patients than in controls under this global coupling scheme.
$3.$  Wilson–Cowan Model with Distinct Coupling Parameters $G_{1}$ and $G_{2}$
@@
Figure 5: Group-averaged best-performing simulation for the SCHZ group using hemispheric-specific coupling parameters $G_{1}=0.80$, $G_{2}=15.00$. (a) Structural Connectivity (SC) matrix, (b) Simulated Functional Connectivity (FC), and (c) Empirical FC. The simulation yielded a correlation of $r=0.47$ and RMSE = 0.20. Panels (a) and (c) correspond to the same subject as in Fig. (3c) and Fig. (3e), respectively.
@@ 4.  Conclusions
In this study, we implemented the Wilson–Cowan model and introduced a hemispheric-specific coupling scheme. We examined both groups and compared structural connectivity network features. In structural connectivity, we found results consistent with previous studies reporting decreased anatomical connectivity strength [9, 15], redistribution of central roles across network regions [15], and topological reorganization in schizophrenia [10]. To further explore the mechanistic basis of these alterations, we evaluated the capacity of a biologically grounded neural mass model based on Wilson–Cowan dynamics [1, 2, 12], in order to reproduce resting-state functional connectivity from empirical cortical structural connectomes in matched control and schizophrenia groups. In our implementation, we differentiated intra-hemispheric and inter-hemispheric anatomical projections via a hemispheric-specific coupling scheme. This modified version systematically outperformed the classic global coupling approach, yielding higher Pearson correlations between simulated and empirical FC across both groups. The improvement was especially pronounced in the schizophrenia group in both models. This may reflect a stronger constraint of functional dynamics by the underlying structural scaffold in schizophrenia, where reduced global integration and increased modularity limit the repertoire of functional configurations. In such conditions, anatomically grounded models may more accurately capture the structure–function mapping [5, 8]. Interestingly, the optimal parameter configuration differed between groups. The SCHZ group achieved a higher maximum correlation of $\langle r\rangle=0.43\pm 0.03$ at $G_{1}=0.8$, $G_{2}=15.0$, whereas the CTRL group reached $\langle r\rangle=0.36\pm 0.05$ at $G_{1}=1.0$, $G_{2}=15.0$. Although both groups showed optimal performance in regimes where inter-hemispheric coupling $G_{2}$ exceeded intra-hemispheric coupling $G_{1}$, the schizophrenia group required a lower $G_{1}$ to reproduce empirical FC. This may suggest a diminished role of intra-hemispheric integration in shaping functional dynamics in schizophrenia, potentially reflecting degraded or less functionally effective intra-hemispheric anatomical pathways [15]. Additionally, the SCHZ model displayed good performance across a broader range of $G_{2}$ values, indicating a degree of flexibility or compensatory reliance on inter-hemispheric interactions. Together, these findings suggest distinct structure–function coupling regimes between groups, with schizophrenia exhibiting both reduced intra-hemispheric integration and a more tolerant configuration space for inter-hemispheric coupling. Overall, these findings highlight the value of integrating structural connectivity data with biologically inspired models to gain mechanistic insights into the altered brain dynamics observed in neuropsychiatric disorders.
Data Availability and Reproducibility
All simulations were implemented in Python 3.11. The implementation relies on the following open-source libraries: NumPy 1.26, SciPy 1.13, pandas 2.2.1, matplotlib 3.8.3, and Numba 0.60 for just-in-time compilation of performance-critical functions. The Wilson–Cowan model and analysis scripts are structured for modularity and extensibility. The entire simulation pipeline can be reproduced by running the provided main scripts (run_global.py and run_hemispheric.py). A requirements.txt file is provided to facilitate environment replication. The simulation outputs (e.g., empirical and simulated functional connectivity matrices, node activity signals, and computed metrics) are stored in a structured directory under results/, which is automatically generated by the code. These outputs include all data used to produce the figures presented in this work. The data that support the findings of this study are openly available in GitHub.